Help please add formula

Saved Horse and Buggy Inc is in a declining industry sales, earnings, and dividends are all shrinking at a rate of 5% per year a. If r: 10% and DVI-$6, what is the value of a share? (Do not round intermediate calculations. Round your answer to 2 decimal places.) b. What price do you forecast for the stock next year? (Do not round intermediate calculations. Round your answer to 2 decimal places.) price c What rote of return should you expect f you buy the stock today and sell t in one year? (Do not round intermediate calculations Enter your answer as a percent rounded to 2 decimal places) rate of return
Answer #1

a) growth rate (g) = - 5%, r = 10%, D1 = $6

Value of share = D1 / (r - g) = $6 / [10% - (-5%)] = $40.00

b) D2 = D1 x (1 + g) = $6 x (1 - 5%) = $5.70

Stock price next year = D2 / (r - g) = $5.70 / [10% - (-5%)] = $38.00

c) rate of return = [ Dividend + (Stock price next price - Value of share today) / Value of share today ] x 100

or, rate of return = [ { $6 + ($38 - $40) } / $40 ] x 100 = 10.00%